A new manufacturing device costs $10,000 to buy, but produces annual benefits of $3250 in reduced production costs. The device will operate for 4 years, at which time it will have no resale value. What is the internal rate of return on this investment?
Answer #1

Initial cost = 10,000

Annual Benefits = 3,250

Life = 4 years

Calculate the internal rate of return.

Calculating the internal rate of return using the trial and error method

Let the rate of interest is 11%. Calculate the PW of the investment at 11%.

PW = -10,000 + 3,250 (P/A, 11%, 4)

PW = -10,000 + 3,250 (3.10245) = 83

The PW is positive. So increase the rate of interest to get negative PW.

Increase the rate of interest to 12% and calculate the PW at 12%.

PW = -10,000 + 3,250 (P/A, 12%, 4)

PW = -10,000 + 3,250 (3.03735) = -129

Using interpolation,

IRR = 11% + [83 – 0 ÷ 83 – (-129)] * 1

IRR = 11.39%

The internal rate of return will be 11.39%.